OpenStudy (anonymous):

Show that y= (2/3)e^x + e^-2x is a solution of the differential equation y' + 2y=2e^x

OpenStudy (bahrom7893):

just take the derivative and substitute

OpenStudy (anonymous):

Is this correct for the derivative y'=2/3e^x+1/e^2?

OpenStudy (bahrom7893):

wait is it: e^(-2x) or [e^(-2)]*x

OpenStudy (anonymous):

its e^(-2x)

OpenStudy (bahrom7893):

okay then the derivative was wrong..

OpenStudy (anonymous):

fml.

OpenStudy (bahrom7893):

y= (2/3)e^x + e^-2x y' = (2/3)e^x + e^(-2x) * (-2)= 2 [ (e^x/3) - e^(-2x) ]

OpenStudy (anonymous):

Lovely Stony Brook university textbook does not have the answer in the back.

OpenStudy (anonymous):

ahhh I forgot to do chain rule didnt I

OpenStudy (bahrom7893):

not only that u just did the differentiation as if it were: e^(-2) * x, but x is in the power.. (e^u)' = e^u (du/dx)

OpenStudy (anonymous):

bloody rusty at derivatives, last semester all we did was integrate hardly any deriving at all

OpenStudy (anonymous):

So im assuming after I plug that in I get 2e^x?

OpenStudy (bahrom7893):

You should get 2e^x once u plug in y' and y where they belong.

OpenStudy (anonymous):

thanks man

OpenStudy (bahrom7893):

np.. I remember how i was killing myself with ODEs last semester haha.. I'm killin myself with lin algebra now..

OpenStudy (anonymous):

Linear algebra will be after differential equations for me. Btw if you wouldnt mind can you attach a file or something to break down plugging in the y and the y'?